OpenStudy (cbecca):

2/3 of the students wore green on St. Patrick`s Day. What fraction of the students did not wear green on St. Patrick`s Day?

OpenStudy (anonymous):

1- 2/3 = ? solve for ?

OpenStudy (cbecca):

sorry did not get it.

OpenStudy (anonymous):

1-2/3 =3/3-2/3 =1/3

OpenStudy (cbecca):

got it, thanks

OpenStudy (anonymous):

np